# Enveloping Ellis semigroups as compactifications of transformations groups

### v2

@@ Abstract.
Thie first author was partially supported by ISF grant 3187/24
@@ 1. Introduction and preliminary remarks
In § 3.2 it is shown that the Roelcke uniformity on the unitary group of a Hilbert space coincides with the uniformity from the family of the stabilizers of points of the unit sphere and is an Ellis equiuniformity (Theorem 3.23). The result provides an alternative proof that the Roelcke compactification of the unitary group is a semitopological group which coincides with the space of all linear operators of norm $\leq 1$ in the weak operator topology [45].
The method for working with Ellis equiuniformities is detailed in § 2.2 and 2.3. In Theorem 2.7 we present both the topological and uniform variants of Ellis’s construction and establish the extremal property of Ellis uniformity. Proposition 2.14 outlines a condition for the coincidence of the Ellis equiuniformity corresponding to the extended action and equiuniformity correponding to the original one. We demonstrate that the map of the poset of equiuniformities on the phase space to the poset of Ellis equiuniformities preserves order. In Proposition 2.16 we establish a condition when the Ellis equiuniformity is equal to the uniformity on a group from a family of small subgroups. Lastly, in Proposition 2.21 we establish the connection between Ellis equiuniformities for the actions on different phase spaces.
@@
All spaces considered are Tychonoff, and we denote a (topological) space as $(X,\tau)$ where $\tau$ is a topology on (a set) $X$. We order topologies: $\tau^{\prime}\geq\tau$ if $\tau\subset\tau^{\prime}$. An abbreviation “nbd” refers to an open neighborhood of a point. The family of all nbds of the unit $e$ in $G$ is denoted $N_{G}(e)$. A (closed) subgroup $H$ of a topological group $G$ is called neutral if for any $O\in N_{G}(e)$ there exists $V\in N_{G}(e)$ such that $VH\subset HO$ [37]. A proper compactification of $X$ is denoted $(bX,b)$ where $bX$ is compacta, $b:X\to bX$ is a dense embedding.
@@ 1.2. Spaces of maps and semigroup structure on $X^{X}$
A uniformity $\mathcal{U}_{X}$ on $X$ is called an equiuniformity [29] if the action of $G\curvearrowright X$ is saturated (any homeomorphism from $G$ is uniformly continuous) and is bounded (for any $u\in\mathcal{U}$ there are $O\in N_{G}(e)$ and $v\in\mathcal{U}$ such that the cover $Ov=\{OV\ |\ V\in v\}\succ u$). In this case $(G,X,\alpha)$ a $G$-Tychonoff space, the action is extended to the continuous action $\tilde{\alpha}:G\times\tilde{X}\to\tilde{X}$ on the completion $(\tilde{X},\tilde{\mathcal{U}}_{X})$ of $(X,\mathcal{U}_{X})$. The extension $\tilde{\mathcal{U}}_{X}$ of $\mathcal{U}_{X}$ to $\tilde{X}$ is an equiuniformity on $\tilde{X}$.
@@
The topology $\tau_{p}$ of pointwise convergence is an admissible group topology since the action is uniformly equicontinuous with respect to the left uniformity $L$ on $G$ [9, Ch. X, § 3, item 5] and $\tau\geq\tau_{p}$. If $O\in N_{(G,\tau)}(e)$, then the set $[e,O]\in\tau_{p}$ and $[e,O]=O$. Thus, $\tau=\tau_{p}$. ∎
@@ 2.1. A proper enveloping Ellis semigroup compactification of a topological group
Let $G$ be a topological group. A proper enveloping Ellis semigroup compactification of $G$ is a pair $(S,s)$, where $S=(S,\bullet)$ is a compact right topological semigroup, $s:G\to S$ is a topological isomorphism of $G$ on a dense subgroup of $S$ and the restriction of the multiplication to $s(G)\times S$ is continuous.
@@
A totally bounded equiuniformity $\mathcal{E}$ on $G$ is an Ellis equiniformity if $\tilde{G}$ is a proper enveloping Ellis semigroup compactification of $G$ ($(\tilde{G},\tilde{\mathcal{E}})$ is the completion of $(G,\mathcal{E})$).
$\mathbb{E}(G)$ is a poset of Ellis equiniformities on $G$.
@@
Not every $G$-compactification $(bG,b)$ of $G$ is a proper enveloping Ellis semigroup compactification of $G$ (equivalently, not every equiniformity on $G$ is an Ellis equiniformity).
@@
The restriction of isomorphism of posets of compactifications of $X$ and totally bounded uniformities on $X$ [14, Ch.8, § 8.4] to the proper enveloping Ellis semigroup compactifications of $G$ and the Ellis equiniformities on $G$ will work if one shows that a continuous map $\psi:S_{2}\to S_{1}$ of the proper enveloping Ellis semigroup compactifications $(S_{1},s_{1})\leq(S_{2},s_{2})$ of $G$ such that $\psi\circ s_{2}=s_{1}$ is a homomorphism.
@@ 2.2. Diagonal action, agreement of algebraic structures with topology
is commutative by item (2) of Proposition 2.6 and $\imath\times\mbox{{\rm id}}$ is a homeomorphism. Continuity of $\alpha_{\Delta X}$ yeilds continuity of $\circ$. (Continuity of the diagonal action $\alpha_{\Delta X}:G\times X^{X}\to X^{X}$ is equivalent to continuity of the multiplication $\circ:\imath(G)\times X^{X}\to X^{X}$.)
@@
($)𝑑$ $\begin{array}[]{ccc}\quad G&\stackrel{{\scriptstyle\imath}}{{\hookrightarrow}}&X^{X}\\ q_{\sigma}\downarrow&&\quad\downarrow\mbox{{\rm pr}}_{\sigma}\\ \quad G/\mbox{{\rm St}}_{\sigma}&\stackrel{{\scriptstyle\varphi_{\sigma}}}{{\longrightarrow}}&X^{\sigma},\end{array}$
where $\varphi_{\sigma}(g\mbox{{\rm St}}_{\sigma})(x)=\alpha(g,x)$, $x\in\sigma$ ($\varphi_{\sigma}$ is correctly defined since if $g^{\prime}=gh$, $h\in\mbox{{\rm St}}_{\sigma}$, then $\alpha(g^{\prime},x)=\alpha(gh,x)=\alpha(g,\alpha(h,x))=\alpha(g,x)$). Considering all spaces as $G$-spaces (with the action of $G$ on $G$ and $G/\mbox{{\rm St}}_{\sigma}$ by multiplication on the left and the diagonal actions (from the action $\alpha$) on products) the maps in the diagram $(d)$ are $G$-maps.
@@
$\mathcal{U}_{X}^{p}|_{\imath(G)}\subset R_{\Sigma_{X}}\ (\mbox{identifing}\ G=\imath(G)).$
@@
($)𝑜𝑝$ $\begin{array}[]{c}\mbox{for any}\ \sigma\in\Sigma_{X}\\ \varphi_{\sigma}:(G/\mbox{{\rm St}}_{\sigma},\mathcal{U}_{G/\mbox{{\rm St}}_{\sigma}})\to((\mbox{{\rm pr}}_{\sigma}\circ\imath)(G),\mathcal{U}_{X}^{\sigma}|_{(\mbox{{\rm pr}}_{\sigma}\circ\imath)(G)})\ \mbox{is a uniform equivalence}{\rm)},\end{array}$
@@
(a) $\mathcal{U}_{X}^{p}|_{\imath(G)}=R_{\Sigma_{X}}\ (\mbox{identifing}\ G=\imath(G))$,
(b) for any $\mathcal{U}_{X}\subset\mathcal{U}^{\prime}_{X}$ (in particular, the maximal equiuniformity on $X$) ${\mathcal{U}^{\prime}}_{X}^{p}|_{\imath(G)}=R_{\Sigma_{X}}\ (\mbox{identifing}\ G=\imath(G))$ (in particular, the map $\varphi_{\sigma}:(G/\mbox{{\rm St}}_{\sigma},\mathcal{U}_{G/\mbox{{\rm St}}_{\sigma}})\to(\mbox{{\rm pr}}_{\sigma}(\imath(G)),{\mathcal{U}^{\prime}}_{X}^{\sigma}|_{\mbox{{\rm pr}}_{\sigma}(\imath(G))})$, $\sigma\in\Sigma_{X}$, is a uniform equivalence).
@@
The restriction of the diagonal action on the invariant subset $\imath(G)$ is continuous. Identificing $G=\imath(G)$ by Proposition 2.6 one has that the action $\cdot$ of $(G,\tau)$ on $(G,\tau_{u})$ is comtinuous. Hence, $\tau\geq\tau_{u}$.
@@ 2.3. Ellis equiuniformity on a group from an equiuniformity on a phase space
()( ⋆ ) for $\begin{array}[]{l}\mbox{for any}\ y\in bX\setminus X\ \mbox{and any}\ \tilde{\rm U}\in\tilde{\mathcal{U}}_{X}\ \mbox{there exists}\ \sigma\in\Sigma_{X}\ \mbox{and}\ \tilde{\rm V}\in\tilde{\mathcal{U}}_{X}\\ \mbox{such that if}\ (\tilde{\alpha}(g,x),\tilde{\alpha}(h,x))\in\tilde{\rm V},x\in\sigma,\ \mbox{then}\ (\tilde{\alpha}(g,y),\tilde{\alpha}(h,y))\in\tilde{\rm U}.\end{array}$
@@
Let $\mathcal{U}_{X}\subset\mathcal{U}^{\prime}_{X}\in\mathbb{BU}(X)$. For the $G$-compatifications $(bX=b_{\mathcal{U}_{X}}X,b=b_{\mathcal{U}_{X}})$ and $(b^{\prime}X=b_{\mathcal{U}^{\prime}_{X}}X,b^{\prime}=b_{\mathcal{U}^{\prime}_{X}})$ of $X$ (completions of $(X,\mathcal{U}_{X})$ and $(X,\mathcal{U}^{\prime}_{X})$ respectively) let a $G$-map $\varphi:b^{\prime}X\to bX$ be such that $\varphi\circ b^{\prime}=b$ (the map of $G$-compactifications). The restriction $\varphi|_{b^{\prime}(X)(=X)}:(X,\mathcal{U}^{\prime}_{X})\to(X,\mathcal{U}_{X})$ is uniformly continuous.
@@
Let $\tilde{\alpha}:G\times bX\to bX$ and $\tilde{\alpha}^{\prime}:G\times b^{\prime}X\to b^{\prime}X$ be the extensions of $\alpha:G\times X\to X$, $X^{\prime}\subset b^{\prime}X$ is any subset the restricition of $\varphi$ to which is a bijection. The Cartesian product $\varphi^{b^{\prime}X}:(b^{\prime}X)^{b^{\prime}X}\to(bX)^{b^{\prime}X}$ of maps $\varphi$ is a $G$-map. The projection $\mbox{{\rm pr}}_{X^{\prime}}:(bX)^{b^{\prime}X}\to(bX)^{X^{\prime}}$ is a $G$-map, the map $\jmath:(bX)^{X^{\prime}}\to(bX)^{bX}$, $\jmath(f)=h$, where $h(t)=f({\varphi^{-1}(t)})$, $t\in bX$ (identification of coordinates) is also a $G$-map. Their composition $\psi=\jmath\circ\mbox{{\rm pr}}_{X^{\prime}}\circ\varphi^{b^{\prime}X}:(b^{\prime}X)^{b^{\prime}X}\to(bX)^{bX}$ is a $G$-map.
@@
()( ⋆ ⋆ ) if $\begin{array}[]{c}\mbox{\rm if for}\ g,h\in G\ (\alpha(g,x),\alpha(h,x))\in{\rm V}\ \mbox{\rm holds for}\ x\in\sigma,\\ \mbox{\rm then there exists}\ \ g^{\prime}\in g\mbox{{\rm St}}_{\sigma}\ \ \mbox{\rm such that}\ \ h\in O_{{\rm U}}g^{\prime},\\ \end{array}$
where $O_{{\rm U}}=\{h\in G\ |\ (\alpha(h,x),x)\in{\rm U},\ x\in X\}$. Then $\mathfrak{E}_{\alpha}(\mathcal{U}_{X})=R_{\Sigma_{X}}$.
@@
By Theorem 2.7 item IV $\tilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{X})\in\mathbb{E}(G)$ us the least equiuniformity $\mathcal{U}$ on $G$ such that the orbit maps $\tilde{\alpha}_{x}:(G,\mathcal{U})\to(bX,\tilde{\mathcal{U}}_{X})$, $x\in bX$, are uniformly continuous. Hence, the orbit maps $\tilde{\alpha}^{\prime}_{y}:(G,\tilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{X}))\to(bY,\tilde{\mathcal{U}}_{Y})$, $y\in bY$, are uniformly continuous. Again appling Theorem 2.7 item IV, one obtains $\tilde{\mathfrak{E}}_{\alpha^{\prime}}(\mathcal{U}_{Y})\subset\tilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{X})$.
@@ 2.4. The maps $\mathfrak{E}_{*}$ and $\tilde{\mathfrak{E}}_{*}$
(a) The projection $\mbox{{\rm pr}}_{e}:(G^{G},\mathcal{U}^{p})\to(G=\{e\}\times G,\mathcal{U})$ is uniformly continuous. Therefore, $\mbox{{\rm pr}}_{e}|_{\imath(G)}:(\imath(G),\mathfrak{E}_{*}(\mathcal{U}))\to(G,\mathcal{U})$ is uniformly continious. Hence, $\mathcal{U}\subset\mathfrak{E}_{*}(\mathcal{U})$. The inclusion $\mathfrak{E}_{*}(\mathcal{U})\subset\tilde{\mathfrak{E}}_{*}(\mathcal{U})$ is proved in Proposition 2.14.
@@
Therefore, for any entorage $\hat{\rm U}=\{(f,g)\in G\times G\ |\ (\imath(f)(t)=b(f)\bullet t,\imath(g)(t)=b(g)\bullet t)\in{\rm U}\}$, where $t\in bG$, ${\rm U}\in\mathcal{U}$, from the subbase of $\tilde{\mathfrak{E}}_{*}(\mathcal{U})$ there exists ${\rm V}\in\mathcal{U}$ such that
@@
and $\tilde{\mathfrak{E}}_{*}(\mathcal{U})=\tilde{\mathfrak{E}}_{*}(\mathcal{U}^{\prime})$. Appling (a) the last statement is obtained. ∎
@@
Since $L\wedge R$ is an equiuniformity for both actions: multiplication on the left and the right, its precomact reflection $\mathcal{U}_{L\wedge R}$ is a totally bounded equiuniformity for the multiplication on the left and the right. Hence, the actions $\cdot_{l}:G\times G\to G$, $\cdot_{l}(g,x)=gx$ and $\cdot_{r}:G\times G\to G$, $\cdot_{r}(g,x)=xg$ extend to the continuous actions on the completion of $(G,\mathcal{U}_{L\wedge R})$.
@@
3. If $G$ is a locally compact group then, the least totally bounded uniformity $\mathcal{U}_{\alpha}$ on $G$ is an Ellis equiniformity and the one-point Alexandroff compactification $\alpha G$ is a proper semitopological semigroup compactification of $G$ [5].
@@
(1)1( 1 ) $\begin{array}[]{ccclc}{\mathfrak{E}}_{\alpha}(\mathcal{U}_{G/H})&\subset&L\wedge R&\subset&R_{\Sigma_{G/H}}\\ \cap&&&&\cap\\ \widetilde{{\mathfrak{E}}}_{\alpha}(\mathcal{U}_{G/H})&&\subset&&R_{\Sigma_{\beta_{G}{G/H}}},\end{array}$
@@
Let $G$ be a subgroup of the permutation group ${\rm S}(X)$ of a discrete (infinite) space $X$. A discrete space $X$ can be considered as a metric space with distance $1$ between distinct points. Then ${\rm S}(X)$ is an isometry group and $G$ is its subgroup. The permutation topology $\tau_{\partial}$ on $G$ is the smallest admissible group topology (and coincides with the topology of pointwise convergence) for the action $\alpha:G\times X\to X$. The base of the nbds of the unit of $(G,\tau_{\partial})$ is formed by clopen subgroups
@@
3.1.1. The maps $\mathfrak{E}_{\alpha}$ and $\tilde{\mathfrak{E}}_{\alpha}$ for an ultransitive action on a discrete space
@@
It remains to check that $f\in b_{r}{\rm S}(X)$ if $f(\infty)=\infty$, $f$ is a bijection on $Y\subset X$ and $f(X\setminus Y)=\infty$, where $Y$ is an arbitraty subset of $X$. Any nbd of $f$ is of the form $\{g\in(\alpha X)^{\alpha X}\ |\ g(x)\in Of(x),\ x\in\sigma\}\cap b_{r}{\rm S}(X)$, $\sigma\in\Sigma_{X}$, where $Of(x)=f(x)$ if $f(x)\neq\infty$, and $Of(x)=\alpha X\setminus\sigma_{x}$, $\sigma_{x}\in\Sigma_{X}$, if $f(x)=\infty$. By finitness of $\sigma$ and $\sigma_{x}$, $x\in\sigma$ there is $h\in{\rm S}(X)$ such that $h(x)=f(x)$ if $x\in\sigma$ and $f(x)\in X$, and $h(x)\not\in\sigma_{x}$, if $x\in\sigma$ and $f(x)=\infty$. Evidently, $h\in\{g\in(\alpha X)^{\alpha X}\ |\ g(x)\in Of(x),\ x\in\sigma\}\cap b_{r}{\rm S}(X)$ and $f\in b_{r}{\rm S}(X)$.

### v3

@@ Abstract.
The study of the first author was carried out with the help of the Center of Integration in Science, Ministry of Aliyah and Integration, Israel, 8123461 and is supported by ISF grant 3187/24.
@@ 1. Introduction and preliminary remarks
In § 3.2 it is shown that the Roelcke uniformity on the unitary group of a Hilbert space coincides with the uniformity from the family of the stabilizers of points of the unit sphere and is an Ellis equiuniformity (Theorem 3.23). The result provides an alternative proof that the Roelcke compactification of the unitary group is a semitopological group which coincides with the space of all linear operators of norm 1$\leq$ 1≤ 1 in the weak operator topology [45].
The method for working with Ellis equiuniformities is detailed in § 2.2 and 2.3. In Theorem 2.7 we present both the topological and uniform variants of Ellis’s construction and establish the extremal property of Ellis uniformity. Proposition 2.14 outlines a condition for the coincidence of the Ellis equiuniformity corresponding to the extended action and equiuniformity corresponding to the original one. We demonstrate that the map of the poset of equiuniformities on the phase space to the poset of Ellis equiuniformities preserves order. In Proposition 2.16 we establish a condition when the Ellis equiuniformity is equal to the uniformity on a group from a family of small subgroups. Lastly, in Proposition 2.21 we establish the connection between Ellis equiuniformities for the actions on different phase spaces.
@@
[…]
@@ 1.2. Spaces of maps and semigroup structure on $X^{X}$
A uniformity $\mathcal{U}_{X}$ on $X$ is called an equiuniformity [29] if the action $G\curvearrowright X$ is saturated (any homeomorphism from $G$ is uniformly continuous) and is bounded (for any $u\in\mathcal{U}$ there are $O\in N_{G}(e)$ and $v\in\mathcal{U}$ such that the cover $Ov=\{OV\ |\ V\in v\}\succ u$). In this case $(G,X,\alpha)$ a $G$-Tychonoff space, the action is extended to the continuous action $\tilde{\alpha}:G\times\tilde{X}\to\tilde{X}$ on the completion $(\tilde{X},\tilde{\mathcal{U}}_{X})$ of $(X,\mathcal{U}_{X})$. The extension $\tilde{\mathcal{U}}_{X}$ of $\mathcal{U}_{X}$ to $\tilde{X}$ is an equiuniformity on $\tilde{X}$.
@@
[…]
@@ 2.1. A proper enveloping Ellis semigroup compactification of a topological group
[…]
@@
A totally bounded equiuniformity $\mathcal{E}$ on $G$ is an Ellis equiuniformity if $\tilde{G}$ is a proper enveloping Ellis semigroup compactification of $G$ ($(\tilde{G},\tilde{\mathcal{E}})$ is the completion of $(G,\mathcal{E})$).
$\mathbb{E}(G)$ is a poset of Ellis equiuniformities on $G$.
@@
[…]
@@
The restriction of isomorphism of posets of compactifications of $X$ and totally bounded uniformities on $X$ [14, Ch.8, § 8.4] to the proper enveloping Ellis semigroup compactifications of $G$ and the Ellis equiuniformities on $G$ will work if one shows that a continuous map $\psi:S_{2}\to S_{1}$ of the proper enveloping Ellis semigroup compactifications $(S_{1},s_{1})\leq(S_{2},s_{2})$ of $G$ such that $\psi\circ s_{2}=s_{1}$ is a homomorphism.
@@ 2.2. Diagonal action, agreement of algebraic structures with topology
is commutative by item (2) of Proposition 2.6 and $\imath\times\mbox{{\rm id}}$ is a homeomorphism. Continuity of $\alpha_{\Delta X}$ yields continuity of $\circ$. (Continuity of the diagonal action $\alpha_{\Delta X}:G\times X^{X}\to X^{X}$ is equivalent to continuity of the multiplication $\circ:\imath(G)\times X^{X}\to X^{X}$.)
@@
[…]
@@
$\mathcal{U}_{X}^{p}|_{\imath(G)}\subset R_{\Sigma_{X}}\ (\mbox{identifying}\ G=\imath(G)).$
@@
[…]
@@
(a) $\mathcal{U}_{X}^{p}|_{\imath(G)}=R_{\Sigma_{X}}\ (\mbox{identifying}\ G=\imath(G))$,
(b) for any $\mathcal{U}_{X}\subset\mathcal{U}^{\prime}_{X}$ (in particular, the maximal equiuniformity on $X$) ${\mathcal{U}^{\prime}}_{X}^{p}|_{\imath(G)}=R_{\Sigma_{X}}\ (\mbox{identifying}\ G=\imath(G))$ (in particular, the map $\varphi_{\sigma}:(G/\mbox{{\rm St}}_{\sigma},\mathcal{U}_{G/\mbox{{\rm St}}_{\sigma}})\to(\mbox{{\rm pr}}_{\sigma}(\imath(G)),{\mathcal{U}^{\prime}}_{X}^{\sigma}|_{\mbox{{\rm pr}}_{\sigma}(\imath(G))})$, $\sigma\in\Sigma_{X}$, is a uniform equivalence).
@@
The restriction of the diagonal action on the invariant subset $\imath(G)$ is continuous. Identifying $G=\imath(G)$ by Proposition 2.6 one has that the action $\cdot$ of $(G,\tau)$ on $(G,\tau_{u})$ is continuous. Hence, $\tau\geq\tau_{u}$.
@@ 2.3. Ellis equiuniformity on a group from an equiuniformity on a phase space
()( ⋆ ) $\begin{array}[]{l}\mbox{for any}\ y\in bX\setminus X\ \mbox{and any}\ \tilde{\rm U}\in\tilde{\mathcal{U}}_{X}\ \mbox{there exists}\ \sigma\in\Sigma_{X}\ \mbox{and}\ \tilde{\rm V}\in\tilde{\mathcal{U}}_{X}\\ \mbox{such that if}\ (\tilde{\alpha}(g,x),\tilde{\alpha}(h,x))\in\tilde{\rm V},x\in\sigma,\ \mbox{then}\ (\tilde{\alpha}(g,y),\tilde{\alpha}(h,y))\in\tilde{\rm U}.\end{array}$
@@
Let $\mathcal{U}_{X}\subset\mathcal{U}^{\prime}_{X}\in\mathbb{BU}(X)$. For the $G$-compactifications $(bX=b_{\mathcal{U}_{X}}X,b=b_{\mathcal{U}_{X}})$ and $(b^{\prime}X=b_{\mathcal{U}^{\prime}_{X}}X,b^{\prime}=b_{\mathcal{U}^{\prime}_{X}})$ of $X$ (completions of $(X,\mathcal{U}_{X})$ and $(X,\mathcal{U}^{\prime}_{X})$ respectively) let a $G$-map $\varphi:b^{\prime}X\to bX$ be such that $\varphi\circ b^{\prime}=b$ (the map of $G$-compactifications). The restriction $\varphi|_{b^{\prime}(X)(=X)}:(X,\mathcal{U}^{\prime}_{X})\to(X,\mathcal{U}_{X})$ is uniformly continuous.
@@
Let $\tilde{\alpha}:G\times bX\to bX$ and $\tilde{\alpha}^{\prime}:G\times b^{\prime}X\to b^{\prime}X$ be the extensions of $\alpha:G\times X\to X$, $X^{\prime}\subset b^{\prime}X$ is any subset the restriction of $\varphi$ to which is a bijection. The Cartesian product $\varphi^{b^{\prime}X}:(b^{\prime}X)^{b^{\prime}X}\to(bX)^{b^{\prime}X}$ of maps $\varphi$ is a $G$-map. The projection $\mbox{{\rm pr}}_{X^{\prime}}:(bX)^{b^{\prime}X}\to(bX)^{X^{\prime}}$ is a $G$-map, the map $\jmath:(bX)^{X^{\prime}}\to(bX)^{bX}$, $\jmath(f)=h$, where $h(t)=f({\varphi^{-1}(t)})$, $t\in bX$ (identification of coordinates) is also a $G$-map. Their composition $\psi=\jmath\circ\mbox{{\rm pr}}_{X^{\prime}}\circ\varphi^{b^{\prime}X}:(b^{\prime}X)^{b^{\prime}X}\to(bX)^{bX}$ is a $G$-map.
@@
()( ⋆ ⋆ ) $\begin{array}[]{c}\mbox{\rm if for}\ g,h\in G\ (\alpha(g,x),\alpha(h,x))\in{\rm V}\ \mbox{\rm holds for}\ x\in\sigma,\\ \mbox{\rm then there exists}\ \ g^{\prime}\in g\mbox{{\rm St}}_{\sigma}\ \ \mbox{\rm such that}\ \ h\in O_{{\rm U}}g^{\prime},\\ \end{array}$
where $O_{{\rm U}}=\{f\in G\ |\ (\alpha(f,x),x)\in{\rm U},\ x\in X\}$. Then $\mathfrak{E}_{\alpha}(\mathcal{U}_{X})=R_{\Sigma_{X}}$.
@@
By Theorem 2.7 item IV $\tilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{X})\in\mathbb{E}(G)$ us the least equiuniformity $\mathcal{U}$ on $G$ such that the orbit maps $\tilde{\alpha}_{x}:(G,\mathcal{U})\to(bX,\tilde{\mathcal{U}}_{X})$, $x\in bX$, are uniformly continuous. Hence, the orbit maps $\tilde{\alpha}^{\prime}_{y}:(G,\tilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{X}))\to(bY,\tilde{\mathcal{U}}_{Y})$, $y\in bY$, are uniformly continuous. Again applying Theorem 2.7 item IV, one obtains $\tilde{\mathfrak{E}}_{\alpha^{\prime}}(\mathcal{U}_{Y})\subset\tilde{\mathfrak{E}}_{\alpha}(\mathcal{U}_{X})$.
@@ 2.4. The maps $\mathfrak{E}_{*}$ and $\tilde{\mathfrak{E}}_{*}$
(a) The projection $\mbox{{\rm pr}}_{e}:(G^{G},\mathcal{U}^{p})\to(G=\{e\}\times G,\mathcal{U})$ is uniformly continuous. Therefore, $\mbox{{\rm pr}}_{e}|_{\imath(G)}:(\imath(G),\mathfrak{E}_{*}(\mathcal{U}))\to(G,\mathcal{U})$ is uniformly continuous. Hence, $\mathcal{U}\subset\mathfrak{E}_{*}(\mathcal{U})$. The inclusion $\mathfrak{E}_{*}(\mathcal{U})\subset\tilde{\mathfrak{E}}_{*}(\mathcal{U})$ is proved in Proposition 2.14.
@@
Therefore, for any entourage $\hat{\rm U}=\{(f,g)\in G\times G\ |\ (\imath(f)(t)=b(f)\bullet t,\imath(g)(t)=b(g)\bullet t)\in{\rm U}\}$, where $t\in bG$, ${\rm U}\in\mathcal{U}$, from the subbase of $\tilde{\mathfrak{E}}_{*}(\mathcal{U})$ there exists ${\rm V}\in\mathcal{U}$ such that
@@
and $\tilde{\mathfrak{E}}_{*}(\mathcal{U})=\tilde{\mathfrak{E}}_{*}(\mathcal{U}^{\prime})$. Applying (a) the last statement is obtained. ∎
@@
Since $L\wedge R$ is an equiuniformity for both actions: multiplication on the left and the right, its precompact reflection $\mathcal{U}_{L\wedge R}$ is a totally bounded equiuniformity for the multiplication on the left and the right. Hence, the actions $\cdot_{l}:G\times G\to G$, $\cdot_{l}(g,x)=gx$ and $\cdot_{r}:G\times G\to G$, $\cdot_{r}(g,x)=xg$ extend to the continuous actions on the completion of $(G,\mathcal{U}_{L\wedge R})$.
@@
3. If $G$ is a locally compact group then, the least totally bounded uniformity $\mathcal{U}_{\alpha}$ on $G$ is an Ellis equiuniformity and the one-point Alexandroff compactification $\alpha G$ is a proper semitopological semigroup compactification of $G$ [5].
@@
(1)( 1 ) $\begin{array}[]{ccclc}{\mathfrak{E}}_{\alpha}(\mathcal{U}_{G/H})&\subset&L\wedge R&\subset&R_{\Sigma_{G/H}}\\ \cap&&&&\cap\\ \widetilde{{\mathfrak{E}}}_{\alpha}(\mathcal{U}_{G/H})&&\subset&&R_{\Sigma_{\beta_{G}{G/H}}},\end{array}$
@@
Let $G$ be a subgroup of the permutation group ${\rm S}(X)$ of a discrete (infinite) space $X$. A discrete space $X$ can be considered as a metric space with distance 111 between distinct points. Then ${\rm S}(X)$ is an isometry group and $G$ is its subgroup. The permutation topology $\tau_{\partial}$ on $G$ is the smallest admissible group topology (and coincides with the topology of pointwise convergence) for the action $\alpha:G\times X\to X$. The base of the nbds of the unit of $(G,\tau_{\partial})$ is formed by clopen subgroups
@@
3.1.1. The maps $\mathfrak{E}_{\alpha}$ and $\tilde{\mathfrak{E}}_{\alpha}$ for an ultratransitive action on a discrete space
@@
It remains to check that $f\in b_{r}{\rm S}(X)$ if $f(\infty)=\infty$, $f$ is a bijection on $Y\subset X$ and $f(X\setminus Y)=\infty$, where $Y$ is an arbitrary subset of $X$. Any nbd of $f$ is of the form $\{g\in(\alpha X)^{\alpha X}\ |\ g(x)\in Of(x),\ x\in\sigma\}\cap b_{r}{\rm S}(X)$, $\sigma\in\Sigma_{X}$, where $Of(x)=f(x)$ if $f(x)\neq\infty$, and $Of(x)=\alpha X\setminus\sigma_{x}$, $\sigma_{x}\in\Sigma_{X}$, if $f(x)=\infty$. By finitness of $\sigma$ and $\sigma_{x}$, $x\in\sigma$ there is $h\in{\rm S}(X)$ such that $h(x)=f(x)$ if $x\in\sigma$ and $f(x)\in X$, and $h(x)\not\in\sigma_{x}$, if $x\in\sigma$ and $f(x)=\infty$. Evidently, $h\in\{g\in(\alpha X)^{\alpha X}\ |\ g(x)\in Of(x),\ x\in\sigma\}\cap b_{r}{\rm S}(X)$ and $f\in b_{r}{\rm S}(X)$.